# Do Students Write Better Post-AI Support?

### v1

@@ 4 Results
4.1 Correlation Between GenAI Literacy and Writing Performance (RQ1)
Spearman’s rank-order correlations were conducted separately by chatbot engagement condition (Passive: ; Proactive: ) to assess the relationship between GenAI literacy and writing rubric scores at both AI-support and AI-removal stages (Figure 5). The analysis revealed significant positive correlations in the passive chatbot condition (Passive group) at the AI-removal stage, indicating that higher GenAI literacy scores were associated with higher academic writing performance after removing AI support. Specifically, GenAI literacy significantly correlated with Insightfulness (), Visual Data Integration (), Organisation and Coherence (), Linguistic Quality (), Critical Thinking (), and Composite Score (). Additionally, a significant positive correlation was found for the Passive group at the AI-support stage only for Linguistic Quality (). No statistically significant correlations between GenAI literacy and writing rubric metrics were observed within the proactive chatbot condition.
@@ 4.2 Comparisons Between Low and High GenAI Literacy Groups (RQ2)
To further explore differences in writing performance based on GenAI literacy levels, participants in the passive chatbot condition () were grouped into low () and high () literacy categories based on median splits of GenAI literacy scores. Mann-Whitney U tests were performed separately for each rubric dimension at both the AI-support and AI-removal stages. Medium to large differences emerged primarily at the AI-removal stage (without AI support), with high GenAI literacy participants consistently scoring higher across several writing dimensions. Specifically, at the AI-removal stage, significant differences were observed for Insightfulness (), with the high literacy group achieving higher scores (Mdn = 2.75, IQR = 1.00) compared to the low literacy group (Mdn = 2.00, IQR = 0.00); Visual Data Integration (), with higher scores for the high literacy group (Mdn = 2.50, IQR = 0.50) versus low literacy group (Mdn = 2.00, IQR = 0.00); and Organisation and Coherence (), again favouring high literacy (Mdn = 2.75, IQR = 0.50) over low literacy participants (Mdn = 2.00, IQR = 0.50). Additionally, a significant difference was found for the overall Composite Score (), with higher scores among high literacy (Mdn = 13.25, IQR = 3.38) than low literacy participants (Mdn = 11.00, IQR = 1.50). At the AI-support stage (with GenAI assistance), only Linguistic Quality revealed a significant difference (), with the high literacy group scoring higher (Mdn = 3.00, IQR = 0.50) compared to the low literacy group (Mdn = 3.00, IQR = 0.00). On the other hand, participants in the proactive chatbot condition () were similarly grouped into low () and high () GenAI literacy categories based on median splits of GenAI literacy scores. However, Mann-Whitney U tests revealed no significant differences between the low and high literacy groups for any of the rubric metrics at either the AI-support or AI-removal stages.
@@ 4.3 Ordinal Logistic Regression Predicting Writing Performance (RQ3)
For each ordinal logistic regression model, proportional odds assumptions were met (Brant test, all ), and multicollinearity diagnostics indicated no problematic multicollinearity (variance inflation factor [VIF] < 5 for all predictors). Results demonstrated that higher GenAI literacy significantly predicted superior writing performance after AI assistance was withdrawn for three rubric criteria: Visual Data Integration, , , , , ; Critical Thinking, , , , , ; and overall writing quality measured through the Composite Score, , , , , . Specifically, each one-unit increase in standardised GenAI literacy was associated with a 14% greater likelihood of achieving higher ratings in Visual Data Integration, a 15% greater likelihood for Critical Thinking, and an 11% greater likelihood to receive higher overall Composite Score. However, GenAI literacy did not significantly predict scores on the Insightfulness, Organisation and Coherence, and Linguistic Quality dimensions. Additionally, participants’ AI-assisted writing scores significantly predicted their subsequent writing performance across all rubric dimensions: Insightfulness, , , , ; Visual Data Integration, , , , ; Organisation and Coherence, , , , ; Linguistic Quality, , , , ; Critical Thinking, , , , ; and overall Composite Score, , , , . Neither domain-specific knowledge (VLAT scores) nor chatbot interaction type (proactive vs. passive) significantly predicted writing performance in any regression models. Thus, the predictive relationship of GenAI literacy with writing performance was robust regardless of participants’ domain-specific knowledge or chatbot intervention type.
@@ 5 Discussion
This study explored the association between GenAI literacy and students’ multimodal academic writing performance in conditions where either passive or proactive GenAI tools supported academic writing. By examining students’ independent writing performance immediately following the withdrawal of GenAI assistance, we sought to understand whether and how GenAI literacy shapes writing performance once external scaffolding is removed and how interactions with chatbots of different styles might moderate these relationships.
@@ 5.1 GenAI Literacy and Independent Writing Performance (RQ1)
For RQ1, our findings indicated a significant positive relationship between students’ GenAI literacy and their academic writing performance at the independent writing stage immediately after GenAI assistance was removed. Specifically, participants with higher GenAI literacy demonstrated stronger performance across multiple rubric dimensions (insightfulness, visual data integration, organisation and coherence, linguistic quality, critical thinking, and composite scores) when writing without AI support. This result provides empirical evidence underscoring the role of GenAI literacy as potentially influential in learners’ capabilities, particularly when performing independently without technological scaffolds. It appears that GenAI literacy is associated with superior performance in conditions where no external AI prompts or support mechanisms are immediately available. This finding is in line with prior theoretical positions suggesting that effective interactions with technology depend not only on the availability of tools but also on users’ understanding of how best to integrate and leverage them effectively 45, 63. Researchers argue that higher levels of literacy related to emerging digital tools lead to better competence, critical engagement with technology-generated output, and stronger independent strategic thinking 28, 30. Our results reinforce this perspective in the context of GenAI-assisted multimodal writing, highlighting that students with higher GenAI literacy appear better prepared, at least immediately following GenAI assistance, to produce high-quality, independent academic writing without continual external support.
@@ 5.2 The Role of Chatbot Interaction Strategies (RQ2)
Interestingly, for RQ2, significant relationships between higher GenAI literacy and independent academic writing performance emerged only within the passive (reactive) chatbot condition. In contrast, the proactive (actively scaffolding) chatbot condition did not demonstrate a significant correlation between GenAI literacy and subsequent independent writing performance. This finding suggests that chatbot interaction strategies can potentially moderate the relationship between GenAI literacy and student outcomes. One plausible explanation is that passive chatbots, which respond only when students actively query them, require greater cognitive and strategic engagement from learners. GenAI-literate students in such conditions may have had opportunities and abilities to formulate effective prompts, critically evaluate the AI-generated assistance, and transfer the insights gained into their independent writing tasks. Conversely, proactively scaffolding chatbots continuously direct and structure learner engagement, perhaps minimising the need for students to exercise or rely upon their own GenAI literacy skills explicitly. This proactive scaffolding likely enabled both high and low GenAI literacy students to allocate their cognitive resources more effectively during task completion, thus levelling performance differences related to GenAI literacy. Students in the proactive condition benefitted from structured support, reducing the cognitive demands associated with engaging directly with GenAI tools, and enabling those with lower initial GenAI literacy to perform comparably to students with higher sophistication. Previous studies suggest scaffolding can enhance learner performance by decreasing unnecessary cognitive load, especially for those less adept with novel technologies 45, 50. Additionally, by providing structured guidance to learners with varying technological familiarity, this scaffolding approach may help mitigate the risk of a digital divide 8. However, it is also possible that the proactive condition limited opportunities for learners to actively practice and further develop their GenAI literacy skills, a consideration that warrants future investigation. Consequently, educators should balance structured support from GenAI chatbots with opportunities for independent inquiry, encouraging active and strategic engagement that leverages learners’ pre-existing technological literacy.
@@ 5.3 Predictive Role of GenAI Literacy in Multimodal Writing (RQ3)
Ordinal logistic regression (RQ3) analyses further substantiated that GenAI literacy significantly predicted students’ independent writing outcomes. Even after controlling for domain-specific content knowledge, initial AI-assisted writing scores, and chatbot interaction conditions, our analyses indicated GenAI literacy uniquely predicted stronger independent writing outcomes in visual data integration, critical thinking, and overall composite scores post-scaffolding removal. By contrast, domain-specific knowledge and chatbot conditions did not significantly predict independent performance outcomes when considered alongside GenAI literacy. These findings reinforce recent arguments positioning digital tool proficiency (especially with GenAI) as a distinct skill set separate from traditional academic knowledge bases 28, 62, 58. Our findings suggest that training and instruction related explicitly to GenAI literacy may effectively support students in independently performing multimodal, complex writing tasks. This unique predictive influence prompts reconsideration of GenAI literacy as a key pedagogical target, warranting purposeful curriculum inclusion rather than simple assumptions of passive skill acquisition.
@@ 5.4 Implications for Research and Practice
Our findings have several implications for educational research related to the integration of GenAI in academic writing contexts. First, the identified relationships between GenAI literacy and independent multimodal writing performance highlight the importance of incorporating systematic instruction and explicit assessment of students’ technology-related literacy skills within contemporary curricula. Researchers emphasise that students do not naturally or uniformly acquire sophisticated literacy skills required to critically engage with or use advanced AI technologies in academic contexts 28, 45. Therefore, educators and curriculum designers should proactively embed training on critical aspects of GenAI literacy, such as effectively formulating prompts, interpreting and evaluating AI-generated outputs, and iteratively refining generated content in complex multimodal academic tasks. Future researchers might also explore longitudinal trajectories of students’ GenAI literacy development and potential transfer effects of such skills across various academic domains and types of academic tasks 58, 30. Additionally, careful studies into the cognitive and metacognitive strategies students employ in interaction with GenAI could yield valuable pedagogical insights for developing effective human-AI collaboration models 67, 10.
At a practical level, our findings underscore the importance of balancing external GenAI scaffolding support with opportunities for independent exploration and inquiry. Specifically, passive or reactive chatbot interventions appear more directly beneficial to learners with high GenAI literacy, as such interaction strategies require greater student initiative, metacognitive monitoring, and strategic engagement with the technology 16. In instructional practice, educators might strategically incorporate AI-assisted tools with fading levels of scaffolding to progressively allow students to take greater control over their learning processes, aligning with established theoretical foundations (e.g., Zone of Proximal Development) and scaffolding frameworks 56, 57. Practitioners should also deliberately consider design characteristics and interaction patterns of instructional chatbots, ensuring that chatbot designs offer sufficient flexibility while fostering productive human-AI interactions and learner autonomy 45, 63, 60. Taken together, these implications highlight the need for intentional, critical, and well-informed use of GenAI tools in educational practice, enabling learners to utilise these powerful tools effectively while continually strengthening their independent scholarly skills and cognitive competencies.

### v2

@@ 4 Results
4.1 Correlation Between GenAI Literacy, Domain Knowledge, and Writing Performance (RQ1)
Visualisation Literacy and Writing Performance
Spearman’s rank-order correlations between participants’ VLAT scores and writing performance across each rubric and composite score, separated by chatbot condition and time point, indicated no statistically significant associations (Figure 5). In the reactive condition (), correlations ranged from to , all , with the largest being Organisation and Coherence at AI-removal stage (). In the proactive condition (), coefficients ranged from to , again none of the coefficients was significant. These results suggest that domain-specific visualisation literacy did not significantly associate multimodal academic writing performance.
GenAI Literacy and Writing Performance
Spearman’s rank-order correlations revealed that GenAI literacy was positively associated with writing performance, particularly in the reactive chatbot condition during the AI-removal stage (Figure 6). Significant correlations were observed with Insightfulness (), Visual Data Integration (), Organisation and Coherence (), Linguistic Quality (), Critical Thinking (), and the Composite Score (). A smaller but significant effect was also found for Linguistic Quality at the AI-support stage (). In contrast, no significant correlations were observed in the proactive condition.
Partial Correlation Analysis Given the moderate overlap between visualisation literacy and GenAI literacy (reactive: ; Proactive: ), we conducted partial correlations controlling for VLAT scores to rule out confounding effects. The results confirmed that the relationships between GenAI literacy and writing performance remained robust even after controlling for visualisation literacy, particularly in the reactive condition during the AI-removal stage. All significant correlations observed in the zero-order analysis maintained statistical significance in the partial correlation analysis. Specifically, in the reactive condition during AI-removal, partial correlations (controlling for VLAT) were: Insightfulness (), Visual Data Integration (), Organisation and Coherence (), Linguistic Quality (), Critical Thinking (), and Composite Score (). Notably, the partial correlation coefficients were similar to the zero-order correlations, with differences ranging from -0.04 to +0.05, suggesting that the observed relationships between GenAI literacy and writing performance are largely independent of visualisation literacy. In contrast, the proactive condition showed no significant partial correlations at either time point, consistent with the zero-order correlation findings (Figure 7).
@@ 4.2 Comparisons Between Low and High GenAI Literacy Groups (RQ2)
The Mann-Whitney U tests examining visualisation literacy group differences revealed no statistically significant differences between high and low visualisation literacy groups in either condition across any of the measured writing performance metrics or time points (all ). In the reactive condition (), participants were divided based on the median VLAT score into low () and high () groups. Mann-Whitney U tests were performed separately for each rubric dimension at both the AI-support and AI-removal stages. While descriptive patterns suggested that high VLAT participants tended to achieve slightly higher median scores on some metrics during the intervention phase (e.g., Insightfulness: Low VLAT Mdn = 2.00, High VLAT Mdn = 2.50; Organization and Coherence: Low VLAT Mdn = 2.00, High VLAT Mdn = 3.00), these differences were not statistically significant. The largest effect size observed was for Organization and Coherence during the post phase (; small effect), but this did not reach conventional significance levels. Similarly, in the proactive condition (low VLAT (), high VLAT (), no significant differences were observed between VLAT groups, with generally small effect sizes and minimal differences in median scores, particularly during the post phase. These findings suggest that within the context of AI-supported multimodal academic writing tasks, visualisation literacy levels did not substantially differentiate writing performance outcomes, regardless of whether participants received reactive or proactive AI support.
Turning to GenAI literacy, participants in the reactive chatbot condition () were grouped via a within-condition median split into low () and high () literacy groups. Medium to large differences emerged primarily at the AI-removal stage (without AI support), with high GenAI literacy participants consistently scoring higher across several writing dimensions. Specifically, at the AI-removal stage, significant differences were observed for Insightfulness (), with the high literacy group achieving higher scores (Mdn = 2.75, IQR = 1.00) compared to the low literacy group (Mdn = 2.00, IQR = 0.00); Visual Data Integration (), with higher scores for the high literacy group (Mdn = 2.50, IQR = 0.50) versus low literacy group (Mdn = 2.00, IQR = 0.00); and Organisation and Coherence (), again favouring high literacy (Mdn = 2.75, IQR = 0.50) over low literacy participants (Mdn = 2.00, IQR = 0.50). Additionally, a significant difference was found for the overall Composite Score (), with higher scores among high literacy (Mdn = 13.25, IQR = 3.38) than low literacy participants (Mdn = 11.00, IQR = 1.50). At the AI-support stage (with GenAI assistance), only Linguistic Quality revealed a significant difference (), with the high literacy group scoring higher (Mdn = 3.00, IQR = 0.50) compared to the low literacy group (Mdn = 3.00, IQR = 0.00). On the other hand, participants in the proactive chatbot condition () were similarly grouped into low () and high () GenAI literacy categories based on median splits of GenAI literacy scores. However, Mann-Whitney U tests revealed no significant differences between the low and high literacy groups for any of the rubric metrics at either the AI-support or AI-removal stages.
@@ 4.3 Ordinal Logistic Regression Predicting Writing Performance (RQ3)
For each ordinal logistic regression model, proportional odds assumptions were met (Brant test, all ), and multicollinearity diagnostics indicated no problematic multicollinearity (variance inflation factor for all predictors). Results demonstrated that higher GenAI literacy significantly predicted superior writing performance after AI assistance was withdrawn for three rubric criteria: Visual Data Integration, , , , , ; Critical Thinking, , , , , ; and overall writing quality measured through the Composite Score, , , , , . Specifically, each one-unit increase in standardised GenAI literacy was associated with a 14% greater likelihood of achieving higher ratings in Visual Data Integration, a 15% greater likelihood for Critical Thinking, and an 11% greater likelihood to receive higher overall Composite Score. However, GenAI literacy did not significantly predict scores on the Insightfulness, Organisation and Coherence, and Linguistic Quality dimensions. Additionally, participants’ AI-assisted writing scores significantly predicted their subsequent writing performance across all rubric dimensions: Insightfulness, , , , ; Visual Data Integration, , , , ; Organisation and Coherence, , , , ; Linguistic Quality, , , , ; Critical Thinking, , , , ; and overall Composite Score, , , , . Neither domain-specific knowledge (VLAT scores) nor chatbot interaction type (proactive vs. reactive) significantly predicted writing performance in any regression models. Thus, the predictive relationship of GenAI literacy with writing performance was robust regardless of participants’ domain-specific knowledge or chatbot intervention type.
4.4 Mediation Analysis of GenAI Literacy and Writing Performance (RQ4)
The bootstrap mediation analysis was conducted to determine if the relationship between GenAI literacy and independent writing performance was explained by students’ performance on the AI-supported task. The results, summarised in Table 2, did not support the hypothesised mediational pathway (GenAI Literacy AI-Support Performance Independent Performance) in either the reactive or proactive condition.
In the reactive condition, the analysis first confirmed a significant total effect, establishing an overall positive relationship between higher GenAI literacy and stronger independent writing performance (Path c: , ). However, the results did not support the proposed mediational pathway, as the indirect effect was negligible and not statistically significant (Indirect Effect = 0.032, 95% CI [-0.022, 0.108]). Crucially, after accounting for this non-significant indirect path, a notable direct effect between GenAI literacy and independent performance persisted (Path c’: ). This finding demonstrates that the overall benefit of GenAI literacy in the reactive condition was largely independent of the proposed mechanism.
Similarly, in the proactive condition, the indirect effect was not statistically significant (Indirect Effect = 0.022, 95% CI [-0.055, 0.102]). Consistent with prior analyses, GenAI literacy demonstrated no significant total or direct effects on writing performance within this condition, emphasising that it was not a determining factor for student outcomes when proactive scaffolding was provided.
| Parameter | Reactive | Proactive |
| Paths | ||
| Path a (X M) | 0.069 | 0.051 |
| Path b (M Y | X) | 0.468 | 0.438 |
| Path c’ (Direct Effect) | 0.140 | 0.007 |
| Path c (Total Effect) | 0.172* | 0.029 |
| Indirect Effect (a b) | ||
| Point Estimate | 0.032 | 0.022 |
| 95% Confidence Interval | [-0.022, 0.108] | [-0.055, 0.102] |
| Significant | No | No |
| Note: . X=GenAI Literacy, M=AI-Support Score, Y=AI-Removal Score. | ||
@@ 5 Discussion
This study explored the association between GenAI literacy and students’ multimodal academic writing performance in conditions where either reactive or proactive GenAI tools supported academic writing. By examining students’ independent writing performance immediately following the withdrawal of GenAI assistance, we sought to understand whether and how GenAI literacy shapes writing performance once external scaffolding is removed and how interactions with chatbots of different styles might moderate these relationships.
@@ 5.1 GenAI Literacy and Independent Writing Performance (RQ1)
For RQ1, our findings indicated a significant positive relationship between students’ GenAI literacy and their academic writing performance at the independent writing stage immediately after GenAI assistance was removed. Specifically, participants with higher GenAI literacy demonstrated stronger performance across multiple rubric dimensions (insightfulness, visual data integration, organisation and coherence, linguistic quality, critical thinking, and composite scores) when writing without AI support. Importantly, these associations were robust when controlling for domain knowledge, and VLAT scores themselves showed no significant associations with writing scores. This pattern indicates that the observed effects are not attributable to differences in domain knowledge but instead reflect a unique skill set tied to GenAI literacy. This result provides empirical evidence underscoring the role of GenAI literacy as potentially influential in learners’ capabilities, particularly when performing independently without technological scaffolds. It appears that GenAI literacy is associated with superior performance in conditions where no external AI prompts or support mechanisms are immediately available. This finding is in line with prior theoretical positions suggesting that effective interactions with technology depend not only on the availability of tools but also on users’ understanding of how best to integrate and leverage them effectively (Nguyen et al., 2024; Yan et al., 2024a). Researchers argue that higher levels of literacy related to emerging digital tools lead to better competence, critical engagement with technology-generated output, and stronger independent strategic thinking (Jin et al., 2024; Kim et al., 2025). Our results reinforce this perspective in the context of GenAI-assisted multimodal writing, highlighting that students with higher GenAI literacy appear better prepared, at least immediately following GenAI assistance, to produce high-quality, independent academic writing without continual external support.
@@ 5.2 The Role of Chatbot Interaction Strategies (RQ2)
Interestingly, for RQ2, significant relationships between higher GenAI literacy and independent academic writing performance emerged only within the reactive (reactive) chatbot condition. In contrast, the proactive (actively scaffolding) chatbot condition did not demonstrate a significant correlation between GenAI literacy and subsequent independent writing performance. In addition, supplementary Mann–Whitney analyses confirmed that these results were not explained by domain knowledge, as the comparisons of high- and low-VLAT groups revealed no significant differences in writing performance across either condition or time point. This strengthens the interpretation that it was the chatbot interaction strategy, rather than differences in subject knowledge, that shaped whether GenAI literacy translated into stronger independent writing performance. This finding suggests that chatbot interaction strategies can potentially moderate the relationship between GenAI literacy and student outcomes.
One plausible explanation is that reactive chatbots, which respond only when students actively query them, require greater cognitive and strategic engagement from learners. GenAI-literate students in such conditions may have had opportunities and abilities to formulate effective prompts, critically evaluate the AI-generated assistance, and transfer the insights gained into their independent writing tasks. Conversely, proactively scaffolding chatbots continuously direct and structure learner engagement, perhaps minimising the need for students to exercise or rely upon their own GenAI literacy skills explicitly. This proactive scaffolding likely enabled both high and low GenAI literacy students to allocate their cognitive resources more effectively during task completion, thus levelling performance differences related to GenAI literacy. Students in the proactive condition benefitted from structured support, reducing the cognitive demands associated with engaging directly with GenAI tools, and enabling those with lower initial GenAI literacy to perform comparably to students with higher sophistication. Previous studies suggest scaffolding can enhance learner performance by decreasing unnecessary cognitive load, especially for those less adept with novel technologies (Nguyen et al., 2024; Sharma and Hannafin, 2007). Additionally, by providing structured guidance to learners with varying technological familiarity, this scaffolding approach may help mitigate the risk of a digital divide (Celik, 2023). However, it is also possible that the proactive condition limited opportunities for learners to actively practice and further develop their GenAI literacy skills, a consideration that warrants future investigation. Consequently, educators should balance structured support from GenAI chatbots with opportunities for independent inquiry, encouraging active and strategic engagement that leverages learners’ pre-existing technological literacy.
@@ 5.3 Predictive Role of GenAI Literacy in Multimodal Writing (RQ3)
Ordinal logistic regression (RQ3) analyses further substantiated that GenAI literacy significantly predicted students’ independent writing outcomes. Even after controlling for domain-specific content knowledge, initial AI-assisted writing scores, and chatbot interaction conditions, our analyses indicated GenAI literacy uniquely predicted stronger independent writing outcomes in visual data integration, critical thinking, and overall composite scores post-scaffolding removal. By contrast, domain-specific knowledge and chatbot conditions did not significantly predict independent performance outcomes when considered alongside GenAI literacy. These findings reinforce recent arguments positioning digital tool proficiency (especially with GenAI) as a distinct skill set separate from traditional academic knowledge bases (Jin et al., 2024; Yan, 2023; Wang, 2025). Our findings suggest that training and instruction related explicitly to GenAI literacy may effectively support students in independently performing multimodal, complex writing tasks. This unique predictive influence prompts reconsideration of GenAI literacy as a key pedagogical target, warranting purposeful curriculum inclusion rather than simple assumptions of reactive skill acquisition.
5.4 Direct Effect of GenAI Literacy (RQ4)
Our fourth research question examined whether the effect of GenAI literacy on independent writing was mediated by in-task success. The mediation pathway was not significant in either condition. This finding challenges the view that the benefits of GenAI literacy are realised mainly through improved tool-assisted performance.
The results instead point to a direct and foundational competence. In the reactive condition the indirect path was not significant, while the direct path was reliable. This suggests that GenAI literacy does not simply improve the AI-supported attempt but provides students with transferable skills they can apply independently. A possible concern is that GenAI literacy may reflect other literacies (Stolpe and Hallström, 2024), such as general writing ability, information literacy, language proficiency, or digital literacy (Yang, 2022). Three pieces of evidence in our data argue against this. First, domain knowledge relevant to the task, measured as visualisation literacy, showed no reliable association with outcomes and did not reduce the effect of GLAT when controlled. Second, the direct rather than indirect path was observed, which is inconsistent with the view that GLAT only enhances the AI-assisted attempt. Third, GLAT predicted performance only when learner agency was required in the reactive chatbot condition. This pattern is consistent with a construct that captures strategic planning, critical evaluation and metacognitive monitoring specific to working with GenAI, rather than general language or domain-specific skills. We therefore interpret GenAI literacy as a distinct, task-general competence that becomes visible when interaction design requires learners to self-initiate and regulate their engagement with GenAI (Yan et al., 2024a).
This interpretation also fits the null effects in the proactive condition. Structured scaffolding reduced the need for students to exercise these strategic capacities and equalised outcomes regardless of GenAI literacy. The influence of GenAI literacy therefore appears less about transfer from an AI-supported task and more about a latent competency that becomes important when learner agency is emphasised.
@@ 5.5 Implications for Research and Practice
The findings highlight the importance of interaction design in shaping how GenAI literacy translates into academic outcomes. The absence of any GLAT effect in the proactive condition indicates that structured scaffolding can equalise performance across learners with different levels of GenAI literacy. In contrast, when support was reactive, students’ own literacy became decisive, with higher-literacy students demonstrating stronger independent writing. This agency gap highlights two complementary levers available to educators and designers: (i) proactive scaffolding could reduce reliance on GenAI literacy during supported work, while (ii) cultivation of GenAI literacy could strengthen students’ capacity to perform independently once external support is removed.
For instructional practice, these results suggest the value of phasing GenAI-supported learning experiences. Educators might strategically integrate AI-assisted tools with fading levels of scaffolding, thereby ensuring students benefit from structured support initially but are progressively encouraged to self-initiate, monitor, and regulate their engagement. Such designs align with established scaffolding frameworks and the Zone of Proximal Development (Vygotsky and Cole, 1978; Van de Pol et al., 2010). Importantly, chatbot design choices, such as whether to respond reactively or scaffold proactively, should be made deliberately, balancing equitable support with opportunities for learner autonomy (Nguyen et al., 2024; Yan et al., 2024a; Xie et al., 2024). At the same time, the direct role of GenAI literacy points to the need for curriculum integration. Students do not uniformly acquire the strategic skills required to effectively prompt, evaluate, and refine AI-generated content (Jin et al., 2024; Nguyen et al., 2024). Embedding instruction on these literacies into academic writing curricula could enhance immediate performance while also prepare learners for transfer to other academic tasks.
Finally, there are important avenues for future research. Longitudinal studies are needed to trace how GenAI literacy develops over time and whether gains transfer across domains and task types (Wang, 2025; Kim et al., 2025). Further work should also investigate the cognitive and metacognitive processes students draw upon when engaging with AI tools (Yang et al., 2025; Cheng et al., 2024), thereby advancing theoretical models of human–AI collaboration in learning.